OpenStudy (anonymous):

How many solutions does the equation have? 6-a=1/3(9a-18) A. one solution B. infinite solutions C. no solution

OpenStudy (anonymous):

A. 6-a=1/3(9a-18) Distribute the 1/3: 6-a=3a-6 add the 6: 12-a=3a add the a: 12=4a Divide by 4: 3=a That means there is only 1 solution.